# Robust Data-Driven Tube-Based Zonotopic Predictive Control with Closed-Loop Guarantees

### v2

@@ II-A Notations
The set of natural numbers, nonnegative integers, real $n$-dimensional space, and real $n\times m$ matrix space are denoted by $\mathbb{N}$, $\mathbb{Z}_{\geq 0}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{n\times m}$. A positive (resp. negative) definite matrix $A$ is denoted by $A>0$ (resp. $A<0$). For a matrix $A$, $\|A\|_{F}$ is its Frobenius norm. For a vector $x$, $\left\lvert x\right\rvert$ is its element-wise absolute value, $\|x\|$ is its 2-norm, and $\|x\|_{P}^{2}=x^{\top}Px$, where $P>0$. The distance of a point $x$ from a set $\mathcal{S}$ is defined as $\|x\|_{\mathcal{S}}=\text{inf}\{\|x-z\|\big{|}z\in\mathcal{S}\}$. For a matrix $A$, $(A)_{j}$ denotes its $j$th column, $A^{\top}$ its transpose, and $A^{\dagger}$ its Moore-Penrose pseudoinverse. $\mathrm{diag}\left(\cdot\right)$ is the diagonal operator making a diagonal matrix by its arguments. $\mathbb{1}_{n}$ and $\mathbb{0}_{n}$ are the n-dimensional vectors with all entries equal to $1$ and 00, respectively. Where no confusion arises, we use $\mathbb{1}$ and $\mathbb{0}$ to denote matrices with the proper dimensions.
@@ II-B Set representation
$\mathcal{Z}_{1}\ominus\mathcal{Z}_{2}$ =$=\{z_{1}\in\mathbb{R}^{n}\,|\,z_{1}\oplus\mathcal{Z}_{2}\subseteq\mathcal{Z}_{1}\},$
@@
A matrix zonotope is presented as $\mathcal{M}=\langle C_{\mathcal{M}},G_{\mathcal{M}}\rangle$, and its definition is analogous to that of a zonotope. It is characterized by a center matrix $C_{\mathcal{M}}\in\mathbb{R}^{n\times m}$ and $\gamma_{\mathcal{M}}\in\mathbb{N}$ generator matrices $G_{\mathcal{M}}=\begin{bmatrix}G_{\mathcal{M}_{1}}&\cdots&G_{\mathcal{M}_{\gamma_{\mathcal{M}}}}\end{bmatrix}\in\mathbb{R}^{n\times(m\gamma_{\mathcal{M}})}$ [19, p. 52]. A matrix zonotope can be over-approximated by an interval matrix $I_{\mathcal{M}}=\mathrm{int}\left(\mathcal{M}\right)=[\underline{\mathcal{M}},\overline{\mathcal{M}}]$, where ¯$\underline{\mathcal{M}}\in\mathbb{R}^{n\times m}$ and ¯$\overline{\mathcal{M}}\in\mathbb{R}^{n\times m}$ denote its lower bound and upper bound, respectively. The Frobenius norm of the interval matrix $I_{\mathcal{M}}$ is defined as $|\!|\!|I_{\mathcal{M}}|\!|\!|_{F}=\||I_{\mathcal{M}_{c}}|+\Delta\|_{F}$ with $I_{\mathcal{M}_{c}}=0.5(\overline{\mathcal{M}}+\underline{\mathcal{M}})$ and $\Delta=0.5(\overline{\mathcal{M}}-\underline{\mathcal{M}})$ [20].
@@ III Problem formulation
where $A\in\mathbb{R}^{n_{x}\times n_{x}}$ and $B\in\mathbb{R}^{n_{x}\times n_{u}}$ are unknown system dynamics matrices, $x(k)\in\mathcal{Z}_{x}=\langle c_{\mathcal{Z}_{x}},G_{\mathcal{Z}_{x}}\rangle\subseteq\mathbb{R}^{n_{x}}$ and $u(k)\in\mathcal{Z}_{u}=\langle c_{\mathcal{Z}_{u}},G_{\mathcal{Z}_{u}}\rangle\subseteq\mathbb{R}^{n_{u}}$ are respectively the state and the control input of the system at time $k\in\mathbb{Z}_{\geq 0}$, and $w(k)$ denotes the noise. Input and state constraints for a controller design are
@@ IV-A Learning Phase
There exists a matrix $K\in\mathbb{R}^{n_{x}\times n_{u}}$ and a positive definite matrix $P=P^{\top}\in\mathbb{R}^{n_{x}\times n_{x}}$ such that, for any chosen nominal model $\begin{bmatrix}\bar{A}&\bar{B}\end{bmatrix}\in\mathcal{M}_{D}$, the inequality $\bar{A}_{K}^{T}P\bar{A}_{K}-P<0$, with $\bar{A}_{K}=\bar{A}+\bar{B}K$, holds.
@@
$\mathcal{Z}_{\epsilon}=\big{\langle}\mathbb{0}_{n_{x}},\mathrm{diag}(|\!|\!|I_{\mathcal{M}_{D}}|\!|\!|_{F}\,\delta/2,\dots,|\!|\!|I_{\mathcal{M}_{D}}|\!|\!|_{F}\,\delta/2)\big{\rangle},$
@@
To complete the proof and show (8), we over-approximate the model mismatch for all $(x,u)\in\mathcal{Z}_{x}\times\mathcal{Z}_{u}$. Since $\mathcal{Z}_{x}{\times}\mathcal{Z}_{u}$ is compact, it follows that all points $(D_{-})_{j}$, $j=1,\dots,T$, are dense in $\mathcal{Z}_{x}{\times}\mathcal{Z}_{u}$. In other words, there exists some $\delta\geq 0$ such that, for any $[x(k)^{\top}\,u(k)^{\top}]^{\top}\in\mathcal{Z}_{x}\times\mathcal{Z}_{u}$, there exists a $(D_{-})_{j}$ for which $\|[x(k)^{\top}\,u(k)^{\top}]^{\top}-(D_{-})_{j}\|\leq\delta$ holds (details in [22]). Here, $\delta$ represents the covering radius for $D_{-}$. Then, for any $[x(k)^{\top}\,u(k)^{\top}]^{\top}\in\mathcal{Z}_{x}\times\mathcal{Z}_{u}$ and $(D_{-})_{j}$, it holds that
@@ IV-B Control phase
Here, $\bar{u}_{t}=\{\bar{u}(t|t),\dots,\bar{u}(t+N-1|t)\}$ and $\bar{x}_{t}=\{\bar{x}(t|t),\dots,\bar{x}(t+N|t)\}$ are the sequences of input and state decision variables, respectively, predicted at time $t$ over the finite horizon $N$. Let $[x_{s}^{\top}\;u_{s}^{\top}]^{\top}$ denote a desired setpoint of the system (2). The cost function in (12a) contains a stage cost $\ell(\bar{x}(k|t),\bar{u}(k|t))=\|\bar{x}(k|t)-x_{s}\|_{Q}^{2}+\|\bar{u}(k|t)-u_{s}\|_{R}^{2}$, where $Q>0$ and $R>0$ are design parameters. Additionally, it contains a terminal cost characterized by the matrix $P>0$, presented as $\ell_{N}(\bar{x}(t+N|t))=\|\bar{x}(t+N|t)-x_{s}\|_{P}^{2}$. Equation (12b) indicates the nominal system dynamics, while (12c) and (12d) represent the tightened input and state constraints to be satisfied. These constraints implicitly incorporate reachability analysis in the optimal problem (12), differently from ZPC [13], where the allowable reachable sets are explicitly computed at each time step. The relationship between the predicted state at current time $\bar{x}(t|t)$ and the measured state $x(t)$ is specified in (12e). Finally, in (12f), $\mathcal{X}_{N}$ is the terminal constraint on the predicted state at time $t+N$.
@@
There exist a matrix $P=P^{\top}>0$, a matrix $K\in\mathbb{R}^{n_{x}\times n_{u}}$, and a set $\mathcal{X}_{N}\subseteq\mathcal{X}\ominus\mathcal{S}$ such that, for all $\bar{x}\in\mathcal{X}_{N}$, the following conditions hold:
@@
$\mathcal{S}(\theta,\kappa)=(1-\theta)^{-1}\bigoplus^{\kappa-1}_{i=0}(\bar{A}_{K}^{\kappa})^{i}\mathcal{Z}_{\phi}$
@@
To compute $K$ and $P$ we follow [16]. After over-approximating $\mathcal{M}_{D}$ by an interval matrix, we determine $K$ and $P$ to satisfy the inequality condition stated in Assumption 4 for its vertices. Due to practical considerations, including computational complexity associated with high dimensional systems, future research will focus on relaxing the hypothesis of a common Lyapunov function stated in Assumption 4.

### v3

@@ II-A Notations
The set of natural numbers, nonnegative integers, real $n$-dimensional space, and real $n\times m$ matrix space are denoted by $\mathbb{N}$, $\mathbb{Z}_{\geq 0}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{n\times m}$. A positive (resp. negative) definite matrix $A$ is denoted by $A>0$ (resp. $A<0$). For a matrix $A$, $\|A\|_{F}$ is its Frobenius norm. For a vector $x$, $\left\lvert x\right\rvert$ is its element-wise absolute value, $\|x\|$ is its 2-norm, and $\|x\|_{P}^{2}=x^{\top}Px$, where $P>0$. The distance of a point $x$ from a set $\mathcal{S}$ is defined as $\|x\|_{\mathcal{S}}=\text{inf}\{\|x-z\|\big{|}z\in\mathcal{S}\}$. For a matrix $A$, $(A)_{j}$ denotes its $j$th column, $A^{\top}$ its transpose, and $A^{\dagger}$ its Moore-Penrose pseudoinverse. $\mathrm{diag}\left(\cdot\right)$ is the diagonal operator making a diagonal matrix by its arguments. $\mathbb{1}_{n}$ and $\mathbb{0}_{n}$ are the n-dimensional vectors with all entries equal to $1$ and 0, respectively. Where no confusion arises, we use $\mathbb{1}$ and $\mathbb{0}$ to denote matrices with the proper dimensions.
@@ II-B Set representation
$\mathcal{Z}_{1}\ominus\mathcal{Z}_{2}$ $=\{z_{1}\in\mathbb{R}^{n}\,|\,z_{1}\oplus\mathcal{Z}_{2}\subseteq\mathcal{Z}_{1}\},$
@@
A matrix zonotope is presented as $\mathcal{M}=\langle C_{\mathcal{M}},G_{\mathcal{M}}\rangle$, and its definition is analogous to that of a zonotope. It is characterized by a center matrix $C_{\mathcal{M}}\in\mathbb{R}^{n\times m}$ and $\gamma_{\mathcal{M}}\in\mathbb{N}$ generator matrices $G_{\mathcal{M}}=\begin{bmatrix}G_{\mathcal{M}_{1}}&\cdots&G_{\mathcal{M}_{\gamma_{\mathcal{M}}}}\end{bmatrix}\in\mathbb{R}^{n\times(m\gamma_{\mathcal{M}})}$ [19, p. 52]. A matrix zonotope can be over-approximated by an interval matrix $I_{\mathcal{M}}=\mathrm{int}\left(\mathcal{M}\right)=[\underline{\mathcal{M}},\overline{\mathcal{M}}]$, where $\underline{\mathcal{M}}\in\mathbb{R}^{n\times m}$ and $\overline{\mathcal{M}}\in\mathbb{R}^{n\times m}$ denote its lower bound and upper bound, respectively. The Frobenius norm of the interval matrix $I_{\mathcal{M}}$ is defined as $|I_{\mathcal{M}}|\!|\!|_{F}=\||I_{\mathcal{M}_{c}}|+\Delta\|_{F}$ with $I_{\mathcal{M}_{c}}=0.5(\overline{\mathcal{M}}+\underline{\mathcal{M}})$ and $\Delta=0.5(\overline{\mathcal{M}}-\underline{\mathcal{M}})$ [20].
@@ III Problem formulation
where $A\in\mathbb{R}^{n_{x}\times n_{x}}$ and $B\in\mathbb{R}^{n_{x}\times n_{u}}$ are unknown system dynamics matrices, $x(k)\in\mathcal{Z}_{x}=\langle c_{\mathcal{Z}_{x}},G_{\mathcal{Z}_{x}}\rangle\subset\mathbb{R}^{n_{x}}$ and $u(k)\in\mathcal{Z}_{u}=\langle c_{\mathcal{Z}_{u}},G_{\mathcal{Z}_{u}}\rangle\subset\mathbb{R}^{n_{u}}$ are respectively the state and the control input of the system at time $k\in\mathbb{Z}_{\geq 0}$, and $w(k)$ denotes the noise. Input and state constraints for a controller design are
@@ IV-A Learning Phase
There exists a matrix $K\in\mathbb{R}^{n_{x}\times n_{u}}$ and a positive definite matrix $P=P^{\top}\in\mathbb{R}^{n_{x}\times n_{x}}$ such that, for any chosen nominal model $\begin{bmatrix}\bar{A}&\bar{B}\end{bmatrix}\in\mathcal{M}_{D}$, and a positive definite matrix $Q_{K}=Q_{K}^{\top}\in\mathbb{R}^{n_{x}\times n_{x}}$, the inequality $\bar{A}_{K}^{T}P\bar{A}_{K}-P<-Q_{K}$, with $\bar{A}_{K}=\bar{A}+\bar{B}K$, holds.
@@
[…]
@@
To complete the proof and show (8), we over-approximate the model mismatch for all $(x,u)\in\mathcal{Z}_{x}\times\mathcal{Z}_{u}$. Since $\mathcal{Z}_{x}{\times}\mathcal{Z}_{u}$ is compact, it follows that all points $(D_{-})_{j}$, $j=1,\dots,T$, are dense in $\mathcal{Z}_{x}{\times}\mathcal{Z}_{u}$. In other words, there exists some $\delta\geq 0$ such that, for any $x(k)^{\top}\,u(k)^{\top}]^{\top}\in\mathcal{Z}_{x}\times\mathcal{Z}_{u}$, there exists a $(D_{-})_{j}$ for which $\|[x(k)^{\top}\,u(k)^{\top}]^{\top}-(D_{-})_{j}\|\leq\delta$ holds (details in [22]). Here, $\delta$ represents the covering radius for $D_{-}$. Then, for any $x(k)^{\top}\,u(k)^{\top}]^{\top}\in\mathcal{Z}_{x}\times\mathcal{Z}_{u}$ and $(D_{-})_{j}$, it holds that
@@ IV-B Control phase
Here, $\bar{u}_{t}=\{\bar{u}(t|t),\dots,\bar{u}(t+N-1|t)\}$ and $\bar{x}_{t}=\{\bar{x}(t|t),\dots,\bar{x}(t+N|t)\}$ are the sequences of input and state decision variables, respectively, predicted at time $t$ over the finite horizon $N$. Let $x_{s}^{\top}\;u_{s}^{\top}]^{\top}$ denote a desired setpoint of the system (2). The cost function in (12a) contains a stage cost $\ell(\bar{x}(k|t),\bar{u}(k|t))=\|\bar{x}(k|t)-x_{s}\|_{Q}^{2}+\|\bar{u}(k|t)-u_{s}\|_{R}^{2}$, where $Q>0$ and $R>0$ are design parameters. Additionally, it contains a terminal cost characterized by the matrix $P>0$, presented as $\ell_{N}(\bar{x}(t+N|t))=\|\bar{x}(t+N|t)-x_{s}\|_{P}^{2}$. Without loss of generality, we let $u_{s}=\mathbb{0}_{n_{u}}$ and $x_{s}=\mathbb{0}_{n_{x}}$. Equation (12b) indicates the nominal system dynamics, while (12c) and (12d) represent the tightened input and state constraints to be satisfied. These constraints implicitly incorporate reachability analysis in the optimal problem (12), differently from ZPC [13], where the allowable reachable sets are explicitly computed at each time step. The relationship between the predicted state at current time $\bar{x}(t|t)$ and the measured state $x(t)$ is specified in (12e). Finally, in (12f), $\mathcal{X}_{N}$ is the terminal constraint on the predicted state at time $t+N$.
@@
There exist a matrix $P=P^{\top}>0$, a matrix $K\in\mathbb{R}^{n_{u}\times n_{x}}$, and a set $\mathcal{X}_{N}\subseteq\mathcal{X}\ominus\mathcal{S}$ such that, for all $\bar{x}\in\mathcal{X}_{N}$, the following conditions hold:
@@
$\mathcal{S}(\theta,\kappa)=(1-\theta)^{-1}\bigoplus^{\kappa-1}_{i=0}\bar{A}_{K}^{i}\mathcal{Z}_{\phi}$
@@
To compute $K$ and $P$ we follow [16]. After over-approximating $\mathcal{M}_{D}$ by an interval matrix, we determine $K$ and $P$ to satisfy the inequality condition stated in Assumption 4, with $Q_{K}=Q+K^{\top}RK$ for its vertices. Due to practical considerations, including computational complexity associated with high dimensional systems, future research will focus on relaxing the hypothesis of a common Lyapunov function stated in Assumption 4.